OpenStudy (anonymous):

Solve for x. (x + 12)(x + 12) = 0 please help !

OpenStudy (anonymous):

(x + 12) (x +12)=0 [(x + 12)]^2 = 0 square rooting both sides; => x + 12 =0 therefore, x = -12

OpenStudy (anonymous):

thank you so much !

OpenStudy (anonymous):

No problem! :)